Use Euclid’s division algorithm to find the HCF of: 
 867 and 225

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

867 and 225
As we know, 867 is greater than 225. Let us apply nowEuclid’s division algorithm on 867, to get,
867 = 225 × 3 + 102
Remainder 102 ≠ 0, therefore taking 225 as divisor andapplying the division lemma method, we get, 225 = 102 × 2 + 51
Again, 51 ≠ 0. Now 102 is the new divisor, so repeating the same step we get,
102 = 51 × 2 + 0
the remainder is now zero, so our procedure stops here. Since, in the last step, the divisor is 51, therefore, HCF (867,225) = HCF(225,102) = HCF(102,51) = 51.
Hence, the HCF of 867 and 225 is 51.

Thanks